OpenStudy (anonymous):

Find general solution of each differential equation. Need help separating the variables? dy/dx = e^x-y

OpenStudy (ribhu):

e^y dy = e^x dx integrate both the sides

OpenStudy (anonymous):

How I'd ypou just separate the x-y onto each side? Is that a property of e?

OpenStudy (anonymous):

*did you

OpenStudy (ribhu):

e^x-y = e^x / e^y

OpenStudy (anonymous):

Ah okay that makes sense. Thanks again. You are the best

OpenStudy (ribhu):

just be my guest and i'll do it all for u